Before iodide was added to our table salt , goiter disease was quite common. you are investigating the occurrence of loiter disease in an isolated community by testing ground water for traces I-. A scientist prepares a blank solution from distilled water and the absorbance was measured to be 0.049. The scientist then prepares a 5.91 mol/L solution of the coloured complex and measures an absorbance of 0.377 at 635 nm in a 1.000 cm cuvette. The scientist takes a ground water sample and prepares it as described above in a 1.000 cm cuvette and measures the absorbance of 0.161. Correct the analyses for the blank and report the concentration of I- in ground water to 2 decimal places in units of umol/L.
Answer #1

Blank solution absorbance = 0.049

Coloured complex absorbance after correction = 0.377-0.049 = 0.328

A = \epsilon lC

A= absorbance = 0.328

\epsilon = molar absorptivity

l = path length = 1 cm

C = 5.91 M

0.328 = \epsilon x 1 cm x 5.91 M

Molar absorptivity, \epsilon = 0.328/5.91 = 0.0555 /M cm

Ground water absorbance after correction = 0.161 - 0.049 = 0.112

A = \epsilon lC

0.112 = 0.0555 /M cm x 1 cm x C

0.112/0.0555 = C

C = 2.02 mol/l

Concentration of iodide in the ground water, C = 2.02 M = 2.02 mol/l = 2020000 x 10-6 M = 2020000.00 uM ( 10-6 = 1micro) (assuming that all the iodide in the sample formed iodine and that form complex).